OpenStudy (anonymous):

catelyn has 2 summer jobs. each week, she works at least 15 hours at a pet store and at least 6 hours as a nanny. She earns $10/hour at the pet store and $8/hour as a nanny. Catelyn wants to work no more than 30 hours and earn at least $250/week. Let x represent the number of hours catelyn works at the pet shop in one week and y represent the number of hours she works as nanny in one week. Write an inequalities that model the four constraints in this situation

OpenStudy (anonymous):

@Hero @robtobey @eric_d

hero (hero):

x = number of hours catelyn works at a pet shop y = number of hours catelyn works as a nanny \(x \ge 15\) \(y \ge 6\) \(10x + 8y \ge 250\) \(x + y \le 30\) Something like that

OpenStudy (anonymous):

ok thanks, i understand now :)